OpenStudy (tori040403):

help will fan & medal! Evaluate. (10 * 4 + 2)(9 + 6\3)

OpenStudy (tori040403):

@SolomonZelman how do i do this again?

OpenStudy (anonymous):

(40+2)(9+2)

OpenStudy (tori040403):

oh thx!

OpenStudy (anonymous):

the answer should be 462

OpenStudy (tori040403):

thx so much!

OpenStudy (anonymous):

no problem. if you need anymore help, you know who to call. :-)

OpenStudy (tori040403):

:-)

OpenStudy (tori040403):

im kinda stuck i forgot which is which (the properties) Which property does the statement illustrate? Column A Column B 1. (p + 3) + q = p + (3 + q) 2. 8 + (9 + 4) = (8 + 9) + 4 3. 4(x – 2) = 4x – 8 4. 7(3 + 5) = 7 • 3 + 7 • 5 A. associative property of addition B. distributive property

OpenStudy (tori040403):

sorry if u see the ?'s

OpenStudy (anonymous):

i answered that question yesterday, let me find the resource i used

OpenStudy (tori040403):

ok thx so much! (yes im from k12)

OpenStudy (anonymous):

do i use A and B that you gave me? or

OpenStudy (tori040403):

yes

OpenStudy (anonymous):

ok. that makes life so much easier.

OpenStudy (tori040403):

where u choose a or b?

OpenStudy (tori040403):

for the numbers

OpenStudy (anonymous):

1. A why do you think? 2. same as 1 why do you think 3 and 4 are b

OpenStudy (anonymous):

i want you to tell me how i got that answer.